### v3

# Simple groups with strong fixed-point properties

Abstract.
We exhibit finitely generated torsion-free groups for which any action on any finite-dimensional CW-complex with finite Betti numbers has a global fixed point.
1. Introduction
In [FFS25], Fournier-Facio and Sun introduced a method of constructing, so called, relative torsion-free Tarski monsters with strong homological control. In this note, we apply their construction to obtain examples of groups with extreme fixed-point properties.
We will say that a group $G$ is a torsion-free Tarski monster if $G$ is a torsion-free infinite simple group all of whose proper subgroups are cyclic.
Theorem 1.1.
There are uncountably many pairwise non-isomorphic finitely generated torsion-free Tarski monsters $G$ such that any action of $G$ on any finite dimensional CW-complex with finite Betti numbers has a global fixed point.
As usual, here and throughout, Betti numbers are taken over $\mathbb{Q}$. Since finite domination, acyclicity and homotopy equivalence to a compact manifold or finite complex all imply finiteness of Betti numbers, 1.1 applies to a broad class of topological spaces commonly studied in geometric topology and group actions.
The study of fixed-point properties for group actions on such spaces has a rich history. A classical result of Smith theory [Smi41] shows that a finite $p$-group acting on a finite-dimensional mod-$p$ acyclic CW-complex must have a fixed point. The first infinite groups with similar fixed-point properties were constructed in [ABJ+09]. Since these constructions rely on torsion, they do not apply to torsion-free groups. More recently, the first torsion-free examples with the property that any action on a finite dimensional contractible CW-complex has a global fixed point were constructed in [FFS25, Cor. C].
It is worth emphasising that in 1.1, neither the assumption of finite-dimensionality nor that of finite Betti numbers can be entirely omitted. For instance, any group acts freely on itself, regarded as a 0-dimensional complex. Moreover, even for simply connected complexes, the omission of either assumption leads to free actions, since any discrete group acts freely on the universal cover of its presentation 2-complex and on the universal cover of its classifying space.
In contrast to 1.1, the next result shows that one can construct torsion-free Tarski monsters of arbitrarily large but finite cohomological dimension, which therefore admit proper actions on finite dimensional contractible CW-complexes. Yet, any action of such a group on a complex of sufficiently smaller dimension with finite Betti numbers must still have a global fixed point. This phenomenon mirrors the behaviour observed in earlier examples of groups with fixed-point properties that depend on the dimension of the space; see, for instance, [ABJ+09, Thm. 1.7], [Ser74, Ser77], [Far09], [BV11], [Bri12], and [Var14].
Theorem 1.2.
For each integer $n\geq 4$, there is a finitely generated torsion-free Tarski monster $G_{n}$ with $\mathrm{cd}(G_{n})=2n$ such that for any $m\leq n-4$, any action of $G_{n}$ on any $m$-dimensional CW-complex with finite Betti numbers has a global fixed point.
The main idea of the proofs of 1.1 and 1.2 is to pit the two spectral sequences associated with an action of the group $G$ on a CW-complex against each other, by comparing their von Neumann dimensions. Since $G$ arises as a quotient of certain even-dimensional hyperbolic arithmetic lattices, its $L^{2}$-Betti numbers can be approximated by those of the lattices, whose spectra have large gaps which we exploit.
Acknowledgments.
We are grateful to Francesco Fournier-Facio, Ian Leary, and Ashot Minasyan for reading an earlier version of this note and for their helpful suggestions. We also thank the referee for a careful reading and for many useful comments.
2. Background
2.1. Homological finiteness properties
A $G$-CW-complex is a CW-complex $X$ equipped with a $G$-action that is compatible with the cell structure, in the sense of Lück [Lüc02, Def. 1.25]. This notion is equivalent to that of an admissible $G$-complex in the terminology of Brown [Bro94, Chap. IX, §10], characterised by the property that any element of $G$ which stabilises a cell $\sigma\subseteq X$ must fix $\sigma$ pointwise. Throughout, we will assume that all $G$-actions are of this form.
We say that $X$ is a free $G$-CW complex if the action $G$ on $X$ is free. By $EG$ we mean a contractible free $G$-CW complex, and by $BG$ we mean the quotient $EG/G$ called a classifying space of $G$. If we do not specify the choice of $EG$, then we simply choose an arbitrary contractible free $G$-CW complex as $EG$. Such a complex always exists and any two choices for $EG$ are $G$-homotopy equivalent [Lüc05]. A group $G$ is said to be of type $F_{n}$ if there is a $BG$ with finite $n$-skeleton and it is of type $F$ if $BG$ is finite.
The cohomological dimension of $G$ can be defined by
$\mathrm{cd}(G)=\sup\{n\in\mathbb{N}\mid H^{n}(G;A)\neq 0\text{ for some }\mathbb{Z}G\text{-module }A\}.$
$2.$ $L^{2}$-Betti numbers
For details, we refer to [Lüc02, Lüc09]. Let $G$ be a discrete group. The group von Neumann algebra of $G$, denoted $\mathcal{N}(G)$, is the set of all bounded linear operators on $\ell^{2}(G)$ that commute with the left regular representation of $G$. The von Neumann dimension $\dim_{\mathcal{N}(G)}$ of $G$ is a function from the set of $\mathcal{N}(G)$-modules to $0,\infty]$ as defined in [Lüc02, Chapter 6], satisfying the following properties.
Proposition 2.3.
Let $G$ be a group. The following properties hold.
  1. (i)
    If $A$ is a finitely generated $\mathcal{N}(G)$-module, then $\dim_{\mathcal{N}(G)}A<\infty$.
  2. (ii)
    If $A$ and $B$ are isomorphic $\mathcal{N}(G)$-modules, then $\dim_{\mathcal{N}(G)}A=\dim_{\mathcal{N}(G)}B$.
  3. (iii)
    If $0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$ is an exact sequence of $\mathcal{N}(G)$-modules, then dim_N(G) B=dim_N(G) A+dim_N(G) C, where the laws of the summation of $0,\infty)$ extends naturally to $0,\infty]$.
  4. (iv)
    Let $\{A_{i}\mid i\in I\}$ be a cofinal system of $\mathcal{N}(G)$-submodules of an $\mathcal{N}(G)$-module $A$, i.e., $A=\bigcup_{i\in I}A_{i}$ and for every two indices $i,j\in I$ there exists an index $k\in I$ such that $A_{i},A_{j}\subset A_{k}$. Then $\dim_{\mathcal{N}(G)}A=\sup\{\dim_{\mathcal{N}(G)}A_{i}\mid i\in I\}$.
Suppose $\theta:K\to G$ is a homomorphism of groups. It induces a ring homomorphism from $\mathbb{Z}K$ to $\mathbb{Z}G$. Since $\mathbb{Z}G$ naturally embeds in $\mathcal{N}(G)$, the ring $\mathcal{N}(G)$ becomes a left $\mathbb{Z}K$-module. So, we can form the chain complex $C_{\ast}(EK)\otimes_{\mathbb{Z}K}\mathcal{N}(G)$ of right $\mathcal{N}(G)$-modules and denote its homology groups by $H_{\ast}(K;\mathcal{N}(G))$ and by
$b^{(2)}_{\ast}(K\xrightarrow{\theta}G):=\dim_{\mathcal{N}(G)}H_{\ast}(K;\mathcal{N}(G)),\;\;b^{(2)}_{\ast}(G):=b^{(2)}_{\ast}(G\xrightarrow{id}G).$
We will only be interested in the special cases where $\theta$ is either a quotient homomorphism or the inclusion homomorphism of a subgroup.
Jaikin-Zapirain and López-Álvarez in [JZLA20] proved Lück’s Approximation Conjecture for virtually locally indicable groups. We will need a quantitative application of this result from [PS25], which is similar to [FFS25, Thm. 2.4].
Corollary 2.4 ([PS25, Cor. 3.4]).
Let $G$ be a finitely generated virtually locally indicable group of type $F_{p+1}$ for some $p\in\mathbb{N}$. Then for every $\delta>0$, there exists a finite subset $\mathcal{F}_{p,\delta}\subset G\smallsetminus\{1\}$ such that if a normal subgroup $N\lhd G$ satisfies $N\cap\mathcal{F}_{p,\delta}=\emptyset$, then for all $n\leq p$,
$|b^{(2)}_{n}(G)-b^{(2)}_{n}(G\twoheadrightarrow G/N)|<\delta.$
3. Proof of Main Results
A key input for the proofs is Theorem 4.1 of [FFS25], of which we state only the parts necessary for our arguments.
Theorem 3.1 (Fournier-Facio–Sun, [FFS25, Thm. 4.1]).
For each positive integer $k$, let $L_{k}$ be a torsion-free non-elementary hyperbolic group, and $B_{k}\subset L_{k}$ be finite subsets. Then there exists a group $M$ such that
  1. (i)
    $M$ is finitely generated torsion-free simple group.
  2. (ii)
    Each proper subgroup of $M$ is cyclic.
  3. (iii)
    There is an epimorphism $\pi_{k}:L_{k}\twoheadrightarrow M$ that is injective on $B_{k}$.
  4. (iv)
    For every $\mathbb{Z}M$-module ( $\mathbb{Z}M$$\mathcal{N}(M)$-bimodule) AA, and $i\geq 3$, there is an isomorphism (of right $\mathcal{N}(M)$-modules) H_i(M; A)≅⨁_k1 H_i(L_k ; A), induced by the epimorphism $\Asterisk_{k\geq 1}L_{k}\twoheadrightarrow M$ whose restriction to each subgroup $L_{k}$ is $\pi_{k}$.
Lemma 3.2.
For each integers $n\geq 1$ and $k\geq 1$, there is a torsion-free virtually locally indicable uniform arithmetic lattice $L_{n,k}$ in $SO(2n,1)$ of type $F$, such that
  1. (i)
    $\mathrm{cd}(L_{n,k})=2n,$
  2. (ii)
    $b^{(2)}_{n}(L_{n,k})>k,$
  3. (iii)
    $b^{(2)}_{j}(L_{n,k})=0,\;\forall j\neq n$.
Proof.
For each $n\geq 1$, let $L_{n}$ be a torsion-free uniform arithmetic lattice of simplest type in $SO(2n,1)$. Since $\mathbb{H}^{2n}/{L_{n}}$ is a closed aspherical manifold, it follows that $L_{n}$ is of type $F$ and $\mathrm{cd}(L_{n})=2n$. By Theorem 2.3 of [JX00], $b^{(2)}_{j}(L_{n})=0$ if $j\neq n$ and $b^{(2)}_{n}(L_{n})=|\chi(L_{n})|$ which is nonzero (see e.g. [PS25, Theorem 7.5]). By Theorem 1.8 of [BW12], each $L_{n}$ is virtually compact special and therefore virtually locally indicable (see e.g. [PS25, Prop. 4.12(iii)]).
Now, for each $k\geq 1$, let $L_{n,k}$ be a finite index subgroup of $L_{n}$ such that $b^{(2)}_{n}(L_{n,k})=|\chi(L_{n,k})|>k$. By [JX00, Thm. 2.3], $b^{(2)}_{j}(L_{n,k})=0$ if $j\neq n$. ∎
Proof of 1.1.
Given an infinite sequence $S=\{n_{l}\;|\;n_{l+1}>n_{l}+l+1\}\subset\mathbb{N}_{\geq 3}$ and $k\geq 1$, let $L_{n_{l},k}$ be hyperbolic lattice groups as in Lemma 3.2. For each pair $(n_{l},k)$, let $B^{\prime}_{l,k}\subset L_{n_{l},k}\smallsetminus\{1\}$ be the finite set given by 2.4 with respect to $p=2n_{l}$ and $\delta=2^{-n_{l}\cdot k}$. We apply Theorem 3.1 with $\{L_{n_{l},k}\}$ for all $l\in\mathbb{N}$ and all $k\geq 1$, and wth the corresponding sets $B_{l,k}:=B^{\prime}_{l,k}\cup\{1\}$, to obtain a finitely generated torsion-free simple group $M_{S}$ such that $\mathrm{cd}(M_{S})=\infty$.
By 3.1(iv), Lemma 3.2 and 2.4,
$\forall j\in S,\;b^{(2)}_{j}(M_{S})$ $=\sum_{k\geq 1,n_{l}\in S}b^{(2)}_{j}(L_{{n_{l}},k}\twoheadrightarrow M_{S})$
$\geq\sum_{k\geq 1}b^{(2)}_{j}(L_{{j},k}\twoheadrightarrow M_{S})$
$\geq\sum_{k\geq 1}(b^{(2)}_{j}(L_{{j},k})-2^{-j\cdot k})=\infty;$
$\forall j\notin S,j\geq 3,\;b^{(2)}_{j}(M_{S})$ $=\sum_{k\geq 1,n_{l}\in S}b^{(2)}_{j}(L_{{n_{l}},k}\twoheadrightarrow M_{S})$
$\leq\sum_{k\geq 1,n_{l}\in S}(b^{(2)}_{j}(L_{{n_{l}},k})+2^{-n_{l}\cdot k})$
$\leq\sum_{k\geq 1,n_{l}\in S}2^{-n_{l}\cdot k}\leq 1.$
This shows that if $S\neq S^{\prime}$, then $\{b^{(2)}_{*}(M_{S})\}\neq\{b^{(2)}_{*}(M_{S^{\prime}})\}$ and hence, $M_{S}\ncong M_{S^{\prime}}$. Since the collection of the above subsets $S$ is uncountable, it follows that, up to isomorphism, the family of the corresponding groups $M_{S}$ is uncountable.
Suppose that for some $S$, the group $M_{S}$ acts on an $m$-dimensional CW-complex $X$ such that $H_{i}(X;\mathbb{Q})$ has finite dimension $b_{i}$ over $\mathbb{Q}$ for each $i$. The action of $M_{S}$ on the complex $X$ induces an action on $H_{i}(X;\mathbb{Q})$ which may be seen as a representation $\rho_{i}:M_{S}\to\mbox{Aut}(H_{i}(X;\mathbb{Q}))=\mathrm{GL}(b_{i},\mathbb{Q})$. By a theorem of Mal’cev [Mal40], any finitely generated subgroup of $\mathrm{GL}(b_{i},\mathbb{Q})$ is residually finite. Since $M_{S}$ is simple, it follows that $\rho_{i}$ is trivial for all $i$, i.e. $M_{S}$ acts trivially on the rational homology of $X$.
Suppose, by a way of contradiction, that the action of $M_{S}$ on $X$ does not have a fixed point.
Next, we recall two standard spectral sequences (see e.g.  [Bro94, §VII.7]) associated to an action of a group on a CW-complex, adapted to our setting. We take coefficients in the ring $\mathcal{N}(M_{S})$, which is a $\mathbb{Z}M_{S}$$\mathcal{N}(M_{S})$-bimodule. This ensures that the spectral sequences naturally have the structure of right $\mathcal{N}(M_{S})$-modules.
Consider the double complex of right $\mathcal{N}(M_{S})$-modules
$P_{*}\otimes_{\mathbb{Z}M_{S}}\displaystyle{(C_{*}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M_{S}))},$
where $P_{*}$ is a projective resolution of $\mathbb{Z}$ over $\mathbb{Z}M_{S}$ and $C_{*}(X;\mathbb{Q})$ is the cellular chain complex of $X$ with rational coefficients. Let $\mbox{Tot}_{*}$ denote the total complex of this double complex. The double complex gives the $E^{1}$-term spectral sequence (see [Bro94, §VII.5 (5.3) p. 169]),
$E_{p,q}^{1}=H_{q}(M_{S};C_{p}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M_{S}))\Longrightarrow H_{p+q}(\mbox{Tot}_{*}),$
Since
$C_{p}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M_{S})\cong\displaystyle{\bigoplus_{\sigma\in\Sigma_{p}}}\mathbb{Q}M_{S}\otimes_{\mathbb{Q}M_{\sigma}}\mathcal{N}(M_{S}),$
as right $\mathcal{N}(M_{S})$-modules, where $X_{p}$ is the collection of all the $p$-cells and $\Sigma_{p}$ denotes a set of representatives of all the $M_{S}$-orbits in $X_{p}$, and $M_{\sigma}$ is the stabiliser of $\sigma$ , the spectral sequence transforms to
$E_{p,q}^{1}=\displaystyle{\bigoplus_{\sigma\in\Sigma_{p}}}H_{q}(M_{\sigma};\mathcal{N}(M_{S}))\Longrightarrow H_{p+q}(\mbox{Tot}_{*}),$
(see [Bro94, §VII.7 (7.7) p. 173]). According to our hypotheses, each stabiliser is a proper subgroup of $M_{S}$ and hence is cyclic. So, by Theorem 6.37 and Theorem 6.54 (7) of [Lüc02], $b^{(2)}_{q}(M_{\sigma}\hookrightarrow M_{S})=0$ for all $q>0$. Therefore, $\dim_{\mathcal{N}(M_{S})}E_{p,q}^{1}=0$ when $p>m$ or $q>0$. This implies that $\mbox{dim}(H_{*}(\mbox{Tot}_{*}))\leq m$, where
$\mbox{dim}(H_{*}(\mbox{Tot}_{*})):=\max\{i\;|\;\dim_{\mathcal{N}(M_{S})}H_{i}(\mbox{Tot}_{*})\neq 0\}.$
The double complex also gives the $E^{2}$-term spectral sequence
$E_{p,q}^{2}=H_{p}(M_{S};H_{q}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M_{S}))\Longrightarrow H_{p+q}(\mbox{Tot}_{*}),$
(see [Bro94, §VII.7 (7.2) p. 172]). Observe that $E_{p,q}^{2}=0$ when $q>m$. Denote by $\mathcal{N}(M_{S})^{b_{q}}$ the direct sum of $b_{q}$ copies of $\mathcal{N}(M_{S})$, equipped with the diagonal $\mathbb{Z}M_{S}$$\mathcal{N}(M_{S})$-bimodule structure. Since $M_{S}$ acts trivially on $H_{*}(X;\mathbb{Q})$,
$H_{q}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M_{S})\cong\mathcal{N}(M_{S})^{b_{q}}$
as $\mathbb{Z}M_{S}$$\mathcal{N}(M_{S})$-bimodules. The spectral sequences then becomes
$E_{p,q}^{2}=H_{p}(M_{S};\mathcal{N}(M_{S}))^{b_{q}}\Longrightarrow H_{p+q}(\mbox{Tot}_{*}).$
[…]
Figure 1. Shaded region $0\leq q\leq m$ illustrates all possible non-zero terms on the $E^{2}$-page and higher pages. The arrow representing the differential $d^{i}$ starting at $(n_{m+1},0)$ remains within this region precisely when $i\leq m+1$, in which case its terminal point does not reach the vertical line at $n_{m}$.
Let $n_{m},n_{m+1}\in S$. Then, $\dim_{\mathcal{N}(M_{S})}E^{2}_{n_{m+1},0}=b_{0}\cdot b^{(2)}_{n_{m+1}}(M_{S})=\infty$ and if $n_{m}<p<n_{m+1}$, then
$\dim_{\mathcal{N}(M_{S})}E^{2}_{p,q}$ $=\dim_{\mathcal{N}(M_{S})}H_{p}(M_{S};\mathcal{N}(M_{S}))^{b_{q}},$
$=b_{q}\cdot b^{(2)}_{p}(M_{S})\leq b_{q}.$
By 2.3(iii), this implies that $\dim_{\mathcal{N}(M_{S})}E^{i}_{p,q}$ is bounded for all $i\geq 2$ whenever $n_{m}<p<n_{m+1}$. Since $E_{p,q}^{2}=0$ for $q>m$ and by construction, $n_{m+1}-n_{m}>m+1$, the target of the differential $d^{i}_{n_{m+1},0}$ has a finite von Neumann dimension for all $i\geq 2$ (see Figure 1). By, again invoking 2.3(iii), for each $i\geq 2$, we obtain
$\dim_{\mathcal{N}(M_{S})}E^{i+1}_{n_{m+1},0}=\dim_{\mathcal{N}(M_{S})}\ker d^{i}_{{n_{m+1}},0}=\infty.$
Since $E_{n_{m+1},0}^{\infty}=E_{n_{m+1},0}^{m+2}$, we obtain $\dim_{\mathcal{N}(M_{S})}E_{n_{m+1},0}^{\infty}=\infty$. Hence, $\mbox{dim}(H_{*}(\mbox{Tot}_{*}))\geq n_{m+1}$ which contradicts our previous estimate. ∎
Proof of Theorem 1.2.
The proof is similar to the proof of Theorem 1.1. So, we only mention the key differences and outline the rest.
Fix $n\geq 4$. For each $k\geq 1$, let $L_{k}:=L_{n,k}$ be hyperbolic lattice groups as in Lemma 3.2 and let $B^{\prime}_{k}\subset L_{k}\smallsetminus\{1\}$ be the finite set given by 2.4 with respect to $p=2n$ and $\delta=2^{-k}$. We apply Theorem 3.1 with $\{L_{k}\}$ and $B_{k}:=B^{\prime}_{k}\cup\{1\}$, to obtain a finitely generated torsion-free simple group $M$ such that $\mathrm{cd}(M)=2n$.
By Theorem 3.1(iv), Lemma 3.2 and 2.4,
$b^{(2)}_{n}(M)$ $=\sum_{k=1}^{\infty}b^{(2)}_{n}(L_{k}\twoheadrightarrow M)$
$\geq\sum_{k=1}^{\infty}(b^{(2)}_{n}(L_{k})-2^{-k})=\infty,$
$b^{(2)}_{j}(M)$ $=\sum_{k=1}^{\infty}b^{(2)}_{j}(L_{k}\twoheadrightarrow M),$
$\leq\sum_{k=1}^{\infty}(b^{(2)}_{j}(L_{k})+2^{-k})\leq 1,\;\forall j\neq n,j\geq 3.$
Suppose $M$ acts on an $m$-dimensional CW-complex $X$ such that $m\leq n-4$ and $H_{i}(X;\mathbb{Q})$ has finite dimension $b_{i}$ over $\mathbb{Q}$ for all $i$. By Mal’cev’s Theorem, any finitely generated subgroup of $\mathrm{GL}(b_{i},\mathbb{Q})$ is residually finite. Thus, the action of $M$ on $X$ induces the trivial action on the rational homology of $X$.
Suppose, by a way of contradiction, that $M$ acts without a global fixed point on $X$. Consider the double complex
$P_{*}\otimes_{\mathbb{Z}M}\displaystyle{(C_{*}(X;\mathbb{Q})\otimes_{\mathbb{Q}}\mathcal{N}(M))},$
where $P_{*}$ is a projective resolution of $\mathbb{Z}$ over $\mathbb{Z}M$. It gives the $E^{1}$-term spectral sequence
$E_{p,q}^{1}=\displaystyle{\bigoplus_{\sigma\in\Sigma_{p}}}H_{q}(M_{\sigma};\mathcal{N}(M))\Longrightarrow H_{p+q}(\mbox{Tot}_{*}).$
According to our hypotheses, each stabiliser is a proper subgroup of $M$ and hence is cyclic. Therefore, $\dim_{\mathcal{N}(M_{S})}E_{p,q}^{1}=0$ when $p>m$ or $q>0$. This implies that $\mbox{dim}(H_{*}(\mbox{Tot}_{*}))\leq m$.
The double complex also gives the $E^{2}$-term spectral sequence
$E_{p,q}^{2}=H_{p}(M;\mathcal{N}(M))^{b_{q}}\Longrightarrow H_{p+q}(\mbox{Tot}_{*}).$
It follows that $\dim_{\mathcal{N}(M)}E^{2}_{n,0}=b_{0}\cdot b^{(2)}_{n}(M)=\infty$ and if $p\neq n,\>p\geq 3$, then
$\dim_{\mathcal{N}(M)}E^{2}_{p,q}$ $=\dim_{\mathcal{N}(M)}H_{p}(M;\mathcal{N}(M))^{b_{q}},$
$=b_{q}\cdot b^{(2)}_{p}(M)\leq b_{q}.$
This implies that if $p\neq n,\>p\geq 3$, then $\dim_{\mathcal{N}(M)}E^{i}_{p,q}$ is bounded for $q\leq m$ and vanishes for $q>m$. Note that if $i\leq m+1$, then $n-i\geq n-m-1\geq 3$, because by assumption $m\leq n-4$. So, $\dim_{\mathcal{N}(M)}E^{i}_{n-i,i-1}$ is bounded for $i\leq m+1$ and vanishes for $i>m+1$. Since $d^{i}_{{n},0}:E^{i}_{n,0}\to E^{i}_{n-i,i-1}$, we obtain
$\dim_{\mathcal{N}(M)}E^{i+1}_{n,0}=\dim_{\mathcal{N}(M)}\ker d^{i}_{{n},0}=\infty.$
Thus, $\dim_{\mathcal{N}(M)}E_{n,0}^{\infty}=\infty$ and hence, $\mbox{dim}(H_{*}(\mbox{Tot}_{*}))\geq n$, which contradicts our previous estimate. ∎